OpenStudy (anonymous):

5. Given: csc θ = −(17/15), where 270° ≤ θ ≤ 360° and cot β = −(3/4) where 90° ≤ β ≤ 180°.

OpenStudy (campbell_st):

the angle is in the 3rd quadrant... and csc = 1/cos find theta, then find \[\theta = \arccos \frac{15}{17}\] then theta is in the 3rd quadrant so its 180 plus the angle you found above. same for the next question 2nd quadrant cot = 1/tan find \[\beta = \arctan \frac{4}{3} \] next subtract the angle found above from 180 to find beta.

OpenStudy (callisto):

I think there are some mistakes there... For the first given condition. First, csc θ = 1/sinθ. Second, 270° ≤ θ ≤ 360° is quadrant IV. So, θ should be an angle in quadrant IV. Actually, the question is incomplete. @jufrando has just posted the given conditions in the question but hasn't posted the whole (or exact) question - what we should solve for the question.

OpenStudy (anonymous):

sorry this is what i got: Evaluate the following, give exact value when possible: sec(225°) sin(-15°) sin(75°) Given: csc θ = −(17/15), where 270° ≤ θ ≤ 360° and cot β = −(3/4) where 90° ≤ β ≤ 180°. Find the exact value of sin(θ + β). Show all the work.

OpenStudy (anonymous):

or thats what it ask me to do

OpenStudy (anonymous):

aah now that is a question \[\sin(\theta+\beta)=\sin(\theta)\cos(\beta)+\sin(\beta)\cos(\theta)\] so you need the following numbers: \(\sin(\theta);\cos(\beta);\sin(\beta);\cos(\theta)\)

OpenStudy (callisto):

csc θ = −(17/15) sin θ = 1/ [−(17/15)] = -15/17 So, adjacent sides \(=\sqrt{17^2 - 15^2} = \sqrt{64} = 8\) cos θ = 8/15 (since it is in quadrant IV, it is positive.)

OpenStudy (anonymous):

\[\csc(\theta)=-\frac{17}{15}\implies \sin(\theta)=-\frac{15}{17}\] so we got one

OpenStudy (callisto):

cot β = −(3/4) tan β = 1/tanβ = 1/ [-(3/4)] = -4/3 So, opposite side = 4 adjacent side = 3 hypotenuse = \(\sqrt{4^2 + 3^3} = \sqrt{25} = 5\) sin β = 4/5 (positive since it is in quad. II) cos β = -3/5 (negative since it is in quad. II)

OpenStudy (callisto):

Special thanks to @satellite73 who gave you the formula you need. Now, plug in the values and get your answer :)

OpenStudy (anonymous):

Oh ok thank you very much for both of you guys :), now i get it :D

OpenStudy (callisto):

Welcome :)